OpenStudy (anonymous):

A snowball melts so that its volume decreases at a rate of 1 cm^3/min. find the rate at which the radius is changing when the diameter is 6 cm.

OpenStudy (anonymous):

V=(4/3)*pi*r^3 (dv/dt)=4*pi*r^2*(dr/dt) (dv/dt)=-1cm^3/sec -1=4*pi*r^2(dr/dt) r=d/2=6/2=3 -1=4*pi*3^2*(dr/dt) (dr/dt)=-1/(36*pi) =-.087266 cm/s